2. Solve the following systems using elimination: -2x + y = 1 (a) , 4x - 2y = -2
Answer #1

2.a) - 2a ty=1 .. ☺ 44-2y=-2 - © from equn ③ x 2 + equn ☺ we get, - 4x +2y=2 42 - 2y = -2 - O - O Then we have infinite inter